# TreeFive.bib

### v1

@@ Abstract.
We prove that a tree with a vertex of degree at least five must be missing a connected partition of some type and therefore its chromatic symmetric function cannot be $e$-positive. We prove that this also holds for a tree with a vertex of degree four as long as it is not adjacent to any leaf. This brings us very close to the conjecture by Dahlberg, She, and van Willigenburg of non-$e$-positivity for all trees with a vertex of degree at least four. We also prove that spiders with four legs cannot have an $e$-positive chromatic symmetric function.
@@
chromatic symmetric function, connected partition, cut vertex, e-positive, proper colouring, spider graph, Stanley–Stembridge conjecture, tree isomorphism conjecture
@@ 1. Introduction
When Stanley [chromsym] defined the chromatic symmetric function $X_{G}$ of a graph $G$, he posed two monumental problems that have garnered tremendous interest. The first is the Stanley–Stembridge conjecture [stanstem], which asserts that $X_{G}$ is $e$-positive whenever $G$ is the incomparability graph of a $(\bm{3}+\bm{1})$-free poset. Many authors proved $e$-positivity for particular subclasses of such graphs [csfvertex, chrombounce3, chromstoe, chromqrook, trilad, chromnsym, qforesttriples] and studied connections to Hessenberg varieties [chromhesssplitting, dothessenberg, hessenberghopf, stanstemhess, chromquasihessenberg] and Hecke algebras [chromhecke, chromcharacters]. The second is the tree isomorphism conjecture, which asserts that trees $T_{1}$ and $T_{2}$ can have $X_{T_{1}}=X_{T_{2}}$ only if they are isomorphic. This is intriguing because all trees with $n$ vertices have the same chromatic polynomial. Many authors proved that the chromatic symmetric function determines certain invariants and therefore distinguishes particular subclasses of trees [propcatchrom, chromtreepolyinv, markedgraphs, propqcatchrom, algochromtrees, catchromiso, distinguishingtrees, chromequal].
Dahlberg, She, and van Willigenburg [chromspostrees] considered a hybrid of these two questions by asking which trees $T$ have that $X_{T}$ is $e$-positive. Many authors [eposclawcon, spider4m22m1, spiderbroom, spiderepos] proved $e$-positivity or non-$e$-positivity for particular subclasses of trees. Dahlberg, She, and van Willigenburg checked the following conjecture for all trees with at most $12$ vertices.
@@
[chromspostrees, Conjecture 6.1] If $T$ is a tree with a vertex of degree $d\geq 4$, then $X_{T}$ is not $e$-positive.
A powerful tool for studying $e$-positivity is the following necessary condition by Wolfgang.
Proposition 1.2.
[conparts, Proposition 1.3.3] A connected graph $G$ must have a connected partition of every type in order for $X_{G}$ to be $e$-positive.
Note that this condition is not sufficient because for example, for every $m\geq 1$ the spider graph $G=S(2m+2,1,1)$ has a connected partition of every type but $X_{G}$ is not $e$-positive. Dahlberg, She, and van Willigenburg proved the following result.
@@
[chromspostrees, Theorem 4.1] If $G$ is an $n$-vertex connected graph with a cut vertex whose deletion produces a graph with $d\geq\log_{2}n+1$ connected components, then $G$ is missing a connected partition of some type and therefore $X_{G}$ is not $e$-positive.
In particular, if $T$ is an $n$-vertex tree with a vertex of degree $d\geq\log_{2}n+1$, then $X_{T}$ is not $e$-positive. Zheng strengthened this result considerably.
@@
[spiderepos, Theorem 3.13] If $G$ is a connected graph with a cut vertex whose deletion produces a graph with at least $6$ connected components, then $G$ is missing a connected partition of some type.
@@
Let $G$ be a connected graph with a cut vertex whose deletion produces a graph with either at least $5$ connected components, or $4$ connected components that each have at least two vertices. Then $G$ is missing a connected partition of some type.
In particular, if $T$ is a tree with a vertex of degree at least $5$, or with a vertex of degree $4$ that is not adjacent to any leaf, then $T$ is missing a connected partition of some type and therefore $X_{T}$ is not $e$-positive.
Note that there exists a tree $T$ with a vertex of degree $4$ that is adjacent to a leaf, and which has a connected partition of every type. Therefore, in this sense, Theorem 1.5 cannot be improved. In fact, we will show in Proposition 3.19 that for every $m\geq 1$ the spider graph $S(6m,6m-2,1,1)$ has a connected partition of every type, so there are infinitely many such examples. This means that to prove non-$e$-positivity of trees with a vertex of degree $4$ in general, we need to know more about the tree to calculate particular coefficients. In the specific case of spiders, we will prove that Conjecture 1.1 holds.
@@
If $S$ is a spider with four legs, then $X_{S}$ is not $e$-positive.
Note that there exist infinite families of spiders with three legs, such as $S(m+1,m,1)$ for $m\geq 1$, that have $e$-positive chromatic symmetric function [chromspostrees, Theorem 3.2].
In the next section, we define chromatic symmetric functions, connected partitions, and spider graphs. In Section 3, we prove Theorem 1.5 by analyzing partial sums of rearrangements of partitions.
@@ 2. Background
A composition is a finite sequence of positive integers $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$, which we call the parts of $\alpha$. We denote by
the set of (proper) partial sums of $\alpha$. A partition is a composition $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ that also satisfies $\lambda_{1}\geq\cdots\geq\lambda_{\ell}$. We write $\lambda\vdash n$ if $\lambda_{1}+\cdots+\lambda_{\ell}=n$. We say that $\alpha$ is a rearrangement of $\lambda$ if every part occurs the same number of times in $\alpha$ and in $\lambda$.
Example 2.1.
The composition $\alpha=(3,4,2,4)$ has $\text{set}(\alpha)=\{3,7,9\}$ and is a rearrangement of the partition $\lambda=(4,4,3,2)\vdash 13$.
Let $G=(V,E)$ be a graph with $n$ vertices. A proper colouring of $G$ is a function $\kappa:V\to\{1,2,3,\ldots\}$ such that whenever vertices $u,v\in V$ are adjacent, we must have $\kappa(u)\neq\kappa(v)$. The chromatic symmetric function of $G$, defined by Stanley [chromsym, Definition 2.1] is the formal power series in infinitely many commuting variables $(x_{1},x_{2},x_{3},\ldots)$ given by
$X_{G}=\sum_{\kappa\text{ proper colouring }}\prod_{v\in V}x_{\kappa(v)}.$
We can write $X_{G}$ uniquely as a linear combination of elementary symmetric functions $e_{\lambda}$, which are indexed by partitions $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})\vdash n$ and defined by
$e_{\lambda}=e_{\lambda_{1}}\cdots e_{\lambda_{\ell}},\text{ where }e_{k}=\sum_{i_{1}\leq\cdots\leq i_{k}}x_{i_{1}}\cdots x_{i_{k}}.$
We say that $X_{G}$ is $e$-positive if all of the coefficients $c_{\lambda}$ in the expansion $X_{G}=\sum_{\lambda\vdash n}c_{\lambda}e_{\lambda}$ are nonnegative. It is also common to say that $G$ itself is $e$-positive. A connected partition of $G$ of type $\lambda\vdash n$ is a collection $\mathcal{S}=\{S_{1},\ldots,S_{\ell}\}$ of pairwise disjoint subsets of $V$ such that
  • $\circ$
    $\bigcup_{i=1}^{\ell}S_{i}=V$,
@@
Proposition 2.2.
[conparts, Proposition 1.3.3] Let $G$ be a connected $e$-positive graph with $n$ vertices. Then for every partition $\lambda\vdash n$, there exists a connected partition of $G$ of type $\lambda$.
A tree is a connected acyclic graph. A leaf of a tree is a vertex of degree one. A cut vertex of a connected graph $G$ is a vertex whose deletion results in a disconnected graph. For a partition $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$, we denote by $S(\lambda)$ the graph consisting of paths of lengths $\lambda_{1},\ldots,\lambda_{\ell}$, called legs, joined at a common vertex. Graphs of the form $S(\lambda)$ are called spiders.
@@
Figure 1 shows some spiders and their chromatic symmetric functions. The spider $S(1,1,1)$ is missing a connected partition of type $(2,2)$ and is indeed not $e$-positive. The spider $S(3,2,1)$ has a connected partition of type $\lambda$ for every $\lambda\vdash 7$ and is $e$-positive. The spider $S(4,1,1)$ has a connected partition of type $\lambda$ for every $\lambda\vdash 7$ but is not $e$-positive, showing that Wolfgang’s necessary condition for $e$-positivity is not sufficient.
Figure 1. Some spiders and their chromatic symmetric functions
$S(1,1,1)$         $S(3,2,1)$         $S(4,1,1)$
$X_{S(1,1,1)}$ $=e_{211}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-2e_{22}}+5e_{31}+4e_{4}$
$X_{S(3,2,1)}$ $=e_{2221}+2e_{3211}+5e_{322}+4e_{331}+12e_{421}+5e_{43}+4e_{511}+13e_{52}+11e_{61}+7e_{7}$
$X_{S(4,1,1)}$ $=e_{2221}+4e_{3211}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-3e_{322}}+10e_{331}+10e_{421}+17e_{43}+4e_{511}+3e_{52}+11e_{61}+7e_{7}$
@@ 3. Proof of main result
We first fix notation that we will use throughout this section. Let $G$ be an $n$-vertex connected graph with a cut vertex $v$ whose deletion produces a graph with at least $3$ connected components, and denote their sizes
@@
We now state our main result using this notation.
@@
Suppose that $c\geq 2$. If any of the following hold, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
@@
    We have $c\geq c_{1}+2$ and $b\geq\frac{c^{2}}{2}$.
In particular, if $c\geq c_{1}+2$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
@@
The condition that $c\geq c_{1}+2$ is precisely that the deletion of $v$ produces either at least $5$ connected components, or $4$ connected components that each have at least two vertices, so Theorem 3.1 implies Theorem 1.5. We cannot remove this hypothesis because the spider graph $S(6,4,1,1)$ has a connected partition of every type $\lambda\vdash 13$.
@@
We cannot remove the hypothesis that $c\geq 2$ because the family of spider graphs $S(m+1,m,1)$ all have a connected partition of every type $\lambda\vdash 2m+3$ and are in fact $e$-positive.
We begin by describing another way to think about connected partitions.
@@
Let $\lambda\vdash n$ be a partition with all parts at least $(c_{1}+1)$. If every rearrangement $\alpha$ of $\lambda$ has a partial sum $s\in\text{set}(\alpha)$ that lies in the interval $I=\{b+1,b+2,\ldots,b+c-1,b+c\}$, then $G$ has no connected partition of type $\lambda$.
@@
Suppose that $G$ has a connected partition $\mathcal{S}$ of type $\lambda$. We will find a rearrangement $\alpha$ of $\lambda$ that has no partial sum in $I$. Let $A,B,C_{1},\ldots,C_{k}$ be the connected components of $G\setminus v$ whose sizes are respectively $a\geq b\geq c_{1}\geq\cdots\geq c_{k}$ and let $S\in\mathcal{S}$ be the subset with $v\in S$. If any $C_{i}\nsubseteq S$, then letting $u\in C_{i}\setminus S$, the subset $S^{\prime}\in\mathcal{S}$ with $u\in S^{\prime}$ would have $|S^{\prime}|\leq c_{i}\leq c_{1}$, which is impossible because every part of $\lambda$ is at least $(c_{1}+1)$. So every $C_{i}\subseteq S$. This means that $\mathcal{S}$ is of the form $\mathcal{S}=\{A_{1},\ldots,A_{i},S,B_{1},\ldots,B_{j}\}$ where $A_{i^{\prime}}\subseteq A$ for $1\leq i^{\prime}\leq i$ and $B_{j^{\prime}}\subseteq B$ for $1\leq j^{\prime}\leq j$. We now take $\alpha$ to be the composition
@@
[…]
We can now prove Part 4 of Theorem 3.1, the case where $b\geq\frac{c^{2}}{2}$.
Proposition 3.6.
If $c\geq c_{1}+2$ and $b\geq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
The idea is to take a partition $\lambda\vdash n$ with all parts equal to $(c-1)$ or $c$. The well-known Frobenius coin problem tells us that every integer $n\geq(c-1)(c-2)$ can be written in the form $n=a_{1}c+a_{2}(c-1)$ for some integers $a_{1},a_{2}\geq 0$. Because $b\geq\frac{c^{2}}{2}$, we have
@@
so we can write $n=a_{1}c+a_{2}(c-1)$ and take $\lambda$ to be $a_{1}$ $c$’s followed by $a_{2}$ $(c-1)$’s. All parts of $\lambda$ are at least $c-1\geq c_{1}+1$. Let $\alpha$ be a rearrangement of $\lambda$. Because $c\geq c_{1}+2\geq 3$, we have $\alpha_{1}\leq c\leq\frac{c^{2}}{2}\leq b$, so there exists some maximal $i$ with $\alpha_{1}+\cdots+\alpha_{i}\leq b$, and then
@@
We now describe another way to ensure that all rearrangements have a partial sum in $I$.
@@
Let $q$ be a positive integer and let $J=\{x,x+1,\ldots,y-1,y\}$, where
@@
Suppose that $x\geq c_{1}+1$. If $\lambda\vdash n$ is a partition with all parts in the interval $J$, then $G$ has no connected partition of type $\lambda$.
@@
Let $x$ and $y$ be positive integers with $x<y$. Then there exists a partition $\lambda\vdash n$ with all parts in the interval $J=\{x,\ldots,y\}$ as long as
@@
For a positive integer $t$, let $tJ=\{tx,tx+1,\ldots,ty-1,ty\}$. First note that every integer $n\in tJ$ has a partition $\lambda\vdash n$ consisting of $t$ integers in $J$. Indeed, for $n=ty$ we can take $\lambda=(y,\ldots,y)$, and for $tx\leq n\leq ty-1$ we can write $n=tq+r$ for some integers $x\leq q\leq y-1$ and $10\leq r\leq t-$ 1, and take $\lambda$ to be $r$ $(q+1)$’s followed by $(t-r)$ $q$’s.
We now see that as $t$ increases, the intervals $tJ$ become closer to each other. Specifically, for $t\geq\left\lceil\frac{x-1}{y-x}\right\rceil\geq\frac{x-1}{y-x}$, we have $ty+1\geq(t+1)x$, which means that as soon as we exit the interval $tJ$, we are already in the interval $(t+1)J$. This means that every integer $n$ satisfying (3.2) is in some interval $tJ$ and has a partition $\lambda\vdash n$ with all parts in $J$. ∎
@@
Proposition 3.10.
If $n\geq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.

### v2

# Graphs missing a connected partition

@@ Abstract.
We prove that a graph with a cut vertex whose deletion produces at least five connected components must be missing a connected partition of some type. We prove that this also holds if there are four connected components that each have at least two vertices. In particular, the chromatic symmetric function of such a graph cannot be $e$-positive. This brings us very close to the conjecture by Dahlberg, She, and van Willigenburg of non-$e$-positivity for all trees with a vertex of degree at least four. We also prove that spiders with four legs cannot have an $e$-positive chromatic symmetric function.
@@
chromatic symmetric function, connected partition, cut vertex, e-positive, graph partition, Stanley–Stembridge conjecture, tree isomorphism conjecture
@@ 1. Introduction
Graph partitioning is a fundamentally important problem in computer science. For example, the max-flow min-cut theorem [7] is a cornerstone result in optimization theory. There are numerous practical applications for finding communities in large networks and graph partitions can allow for parallelization of algorithms to reduce computational complexity. Surveys of graph partitioning problems and algorithms can be found in [1, 3, 11, 12].
Let $G=(V,E)$ be an $n$-vertex graph and let $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ be an integer partition of $n$. A connected partition of $G$ of type $\lambda$ is a partition of the vertices $V$ into connected subsets of sizes $\lambda_{1},\ldots,\lambda_{\ell}$. A more precise definition and examples will be given in Section 2.
Many authors studied problems of finding connected partitions that satisfy certain additional constraints or optimizations [2, 4, 8, 9, 10, 13]. In this paper, we consider the problem of finding an integer partition $\lambda$ for which $G$ is missing a connected partition of type $\lambda$. One application of this is to non-$e$-positivity of chromatic symmetric functions. Stanley [14] defined the chromatic symmetric function of a graph $G=(V,E)$ to be
$X_{G}=\sum_{\begin{subarray}{c}\kappa:V\to\{1,2,3,\ldots\}\\ \text{ if }ij\in E,\text{ then }\kappa(i)\neq\kappa(j)\end{subarray}}\prod_{v\in V}x_{\kappa(v)}$
and we say that $X_{G}$ is $e$-positive if it can be written as a nonnegative linear combination in the basis of elementary symmetric functions, defined by
$e_{\lambda}=e_{\lambda_{1}}\cdots e_{\lambda_{\ell}},\text{ where }e_{k}=\sum_{i_{1}<\cdots<i_{k}}x_{i_{1}}\cdots x_{i_{k}}.$
It is also common to say that $G$ itself is $e$-positive. Wolfgang found a powerful necessary condition for $e$-positivity.
Lemma 1.1.
[18, Proposition 1.3.3] If a connected graph $G$ is missing a connected partition of some type $\lambda\vdash n$, then it is not $e$-positive.
Dahlberg, She, and van Willigenburg [6] were interested in determining which trees are $e$-positive. Many authors [5, 6, 16, 17, 19] proved $e$-positivity or non-$e$-positivity for particular subclasses of trees. Dahlberg, She, and van Willigenburg checked the following conjecture for all trees with at most $12$ vertices.
@@
[6, Conjecture 6.1] If $T$ is a tree with a vertex of degree at least $4$, then $T$ is not $e$-positive.
They proved the following result.
@@
[6, Theorem 4.1] If $G$ is an $n$-vertex graph with a cut vertex whose deletion produces $d\geq\log_{2}n+1$ connected components, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
In particular, if $T$ is an $n$-vertex tree with a vertex of degree $d\geq\log_{2}n+1$, then $T$ is not $e$-positive. Zheng strengthened this result considerably.
@@
[19, Theorem 3.13] If $G$ is a graph with a cut vertex whose deletion produces at least $6$ connected components, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
@@
Let $G$ be a graph with a cut vertex whose deletion produces either at least $5$ connected components, or $4$ connected components that each have at least two vertices. Then $G$ is missing a connected partition of some type $\lambda\vdash n$.
In particular, if $T$ is a tree with a vertex of degree at least $5$, or with a vertex of degree $4$ that is not adjacent to any leaf, then $T$ is not $e$-positive.
We will show in Proposition 3.18 that there exists an infinite family of trees with a vertex of degree $4$ that is adjacent to a leaf, and which have a connected partition of every type $\lambda\vdash n$. Therefore, in this sense, Theorem 1.5 cannot be improved. This means that to prove non-$e$-positivity of trees with a vertex of degree $4$ in general, we need to know more about the tree to calculate particular coefficients. In the specific case of spiders, which are graphs $S(\mu)$ formed by joining paths of lengths $\mu_{1},\ldots,\mu_{\ell}$, the legs of the spider, at a single vertex, we will prove that Conjecture 1.2 holds.
@@
Spiders with four legs are not $e$-positive.
@@ 2. Background
[…]
$S(6,4,1,1)=$
A graph $G=(V,E)$ consists of a set $V$ of vertices and a set $E$ of edges, which are sets of two vertices. In particular, we do not allow loops, which are edges joining a vertex to itself, or parallel edges, which are edges joining the same pair of vertices. We will always use $n$ to denote the order of $G$, which is the number of vertices. For a subset $S\subseteq V$, the induced subgraph is $G[S]=(S,\{e=vw\in E:\ v,w\in S\})$.
We say that $G$ is connected if for every $v,w\in V$, there is a path from $v$ to $w$, which is a sequence of vertices $(v=v_{1},v_{2}\ldots,v_{k-1},v_{k}=w)$ with every $v_{i}v_{i+1}\in E$. A connected component of $G$ is a maximal subset $S\subseteq V$ such that $G[S]$ is connected. A vertex $v\in V$ of a connected graph is a cut vertex if the graph $G\setminus v$ obtained by deleting $v$ and all incident edges is not connected.
A tree $T=(V,E)$ is a minimally connected graph, meaning that for every $e\in E$, the graph $(V,E\setminus\{e\})$ is not connected. The degree of a vertex $v$ is the number of edges incident to it and a leaf of a tree is a vertex of degree one. Note that in a tree, every vertex is a leaf or a cut vertex. If $v$ is a cut vertex of degree $d$, then $T\setminus v$ has $d$ connected components.
A composition of $n$ is a sequence of positive integers $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ with sum $n$. An (integer) partition of $n$ is a weakly decreasing composition $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ of $n$. These integers $\lambda_{i}$ are called the parts of $\lambda$. We write $\lambda\vdash n$ to mean that $\lambda$ is an integer partition of $n$. We now define the main object of study in this paper.
Definition 2.1.
Let $G=(V,E)$ be an $n$-vertex graph and let $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})\vdash n$. A connected partition of $G$ of type $\lambda$ is a set partition $\mathcal{S}=\{S_{1},\ldots,S_{\ell}\}$ of $V$ so that
@@
We say that $G$ is missing a connected partition of type $\lambda$ if no such $\mathcal{S}$ exists.
@@
The spider graph $G=S(6,4,1,1)$ is shown in Figure 1. We have indicated a connected partition of $G$ of type $\lambda=(5,4,3,1)\vdash 13$. We will see in Proposition 3.18 that $G$ has a connected partition of every type $\lambda\vdash 13$. But by Theorem 1.6, $G$ still fails to be $e$-positive. This shows that Wolfgang’s necessary condition for $e$-positivity is not sufficient.
Example 2.3.
If $n$ is even, a connected partition of $G=(V,E)$ of type $\lambda=(2,2,\ldots,2,2)$ is equivalent to a perfect matching of $G$, which is a subset $U\subseteq E$ such that every vertex is in exactly one edge of $U$.
@@ 3. Proof of main result
[…]
We first fix notation that we will use throughout this section.
Notation 3.1.
Let $G$ be an $n$-vertex connected graph with a cut vertex $v$ whose deletion produces at least $3$ connected components, and denote their orders
@@
We will always use $\lambda$ to denote an integer partition of $n$.
We now restate our main result using this notation.
@@
Let $G$, $a$, $b$, $c$, $c_{1}$, and $n$ be as in Notation 3.1. Suppose that $c\geq 2$. If any of the following hold, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
@@
    We have $b\geq\frac{c^{2}}{2}$ and $c\geq c_{1}+2$.
In particular, if $c\geq c_{1}+2$, then no matter which range $b$ lies in, we can conclude that $G$ is missing a connected partition of some type $\lambda\vdash n$.
The proof of Theorem 3.2 is structured as follows. We begin by reformulating connected partitions of type $\lambda$ in terms of partial sums of rearrangements of $\lambda$ (Lemma 3.7), which we use to prove Part (4). We then describe a strategy for choosing the parts of $\lambda$ (Lemma 3.9), which will prove Parts (1) and (2). To prove Part (3), we will need a technical result to establish an inequality (Lemma 3.17). Our estimates will only be valid if $c$ is sufficiently large, namely $c\geq 500$, but we can check the cases with $2\leq c\leq 499$ by computer (Lemma 3.15). Sage code used to check these cases is provided in the Appendix.
@@
The condition that $c\geq c_{1}+2$ is precisely that the deletion of $v$ produces either at least $5$ connected components, or $4$ connected components that each have at least two vertices, so Theorem 3.2 implies Theorem 1.5. We cannot remove this hypothesis because the spider graph $S(6,4,1,1)$ from Figure 1 has a connected partition of every type $\lambda\vdash 13$. Incidentally, it will follow from Theorem 1.6 that the graph $S(6,4,1,1)$ is not $e$-positive.
@@
We cannot remove the hypothesis that $c\geq 2$ because the family of spider graphs $S(m+1,m,1)$ all have a connected partition of every type $\lambda\vdash 2m+3$ and are in fact $e$-positive [5, Theorem 3.2].
We begin by describing another way to think about connected partitions. A composition $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ is a rearrangement of $\lambda$ if each positive integer occurs the same number of times in $\alpha$ and in $\lambda$. We denote by
the set of (proper) partial sums of $\alpha$.
Example 3.6.
The composition $\alpha=(3,1,5,4)$ is a rearrangement of $\lambda=(5,4,3,1)$ and has set of partial sums $\text{set}(\alpha)=\{3,4,9\}$.
@@
Let $G$, $a$, $b$, $c$, $c_{1}$, $n$, and $v$ be as in Notation 3.1. Let $\lambda$ be a partition of $n$ with all parts at least $(c_{1}+1)$. If every rearrangement $\alpha$ of $\lambda$ has a partial sum $s\in\text{set}(\alpha)$ that lies in the interval $I=\{b+1,\ldots,b+c\}$, then $G$ is missing a connected partition of type $\lambda$.
@@
Suppose that $G$ has a connected partition $\mathcal{S}$ of type $\lambda$. We will find a rearrangement $\alpha$ of $\lambda$ that has no partial sum in $I$. Let $A,B,C_{1},\ldots,C_{k}$ be the connected components of $G\setminus v$ whose orders are respectively $a\geq b\geq c_{1}\geq\cdots\geq c_{k}$ and let $S\in\mathcal{S}$ be the subset with $v\in S$. If any $C_{i}\nsubseteq S$, then letting $u\in C_{i}\setminus S$, the subset $S^{\prime}\in\mathcal{S}$ with $u\in S^{\prime}$ would have $|S^{\prime}|\leq c_{i}\leq c_{1}$, which is impossible because every part of $\lambda$ is at least $(c_{1}+1)$. So every $C_{i}\subseteq S$. This means that $\mathcal{S}$ is of the form $\mathcal{S}=\{A_{1},\ldots,A_{i},S,B_{1},\ldots,B_{j}\}$ where $A_{i^{\prime}}\subseteq A$ for $1\leq i^{\prime}\leq i$ and $B_{j^{\prime}}\subseteq B$ for $1\leq j^{\prime}\leq j$. We now take $\alpha$ to be the composition
@@
We also use the following result, the two-coin solution to the Frobenius coin problem, which gives a sufficient condition to find a partition $\lambda$ of $n$ with specified parts.
Theorem 3.8.
[15] Let $x$ and $y$ be positive integers with $\gcd(x,y)=1$. Then every integer $n\geq(x-1)(y-1)$ can be written in the form $n=a_{1}x+a_{2}y$ for some integers $a_{1},a_{2}\geq 0$.
We can now prove Part (4) of Theorem 3.2, the case where $c\geq c_{1}+2$ and $b\geq\frac{c^{2}}{2}$.
Proof of Part (4) of Theorem 3.2. .
The idea is to take a partition $\lambda$ of $n$ with parts $c$ or (c1)(c-1). Because $b\geq\frac{c^{2}}{2}$, we have
@@
so by Theorem 3.8, we can write $n=a_{1}c+a_{2}(c-1)$ and take $\lambda$ to be $a_{1}$ $c$’s followed by $a_{2}$ (c1)(c-1)’s. All parts of $\lambda$ are at least $c-1\geq c_{1}+1$. Let $\alpha$ be a rearrangement of $\lambda$. Because $c\geq c_{1}+2\geq 3$, we have $\alpha_{1}\leq c\leq\frac{c^{2}}{2}\leq b$, so there exists some maximal $i$ with $\alpha_{1}+\cdots+\alpha_{i}\leq b$, and then the partial sum $\alpha_{1}+\cdots+\alpha_{i}+\alpha_{i+1}$ satisfies
@@
We now describe another way to ensure that all rearrangements of $\lambda$ have a partial sum in the interval $I=\{b+1,\ldots,b+c\}$.
@@
Let $G$, $a$, $b$, $c$, $c_{1}$, and $n$ be as in Notation 3.1. Let $q$ be a positive integer and let $J=\{x,x+1,\ldots,y-1,y\}$, where
@@
Suppose that $x\geq c_{1}+1$. If $\lambda$ is a partition of $n$ with all parts in the interval $J$, then $G$ is missing a connected partition of type $\lambda$.
@@
Let $x$ and $y$ be positive integers with $x<y$. Then there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{x,\ldots,y\}$ as long as
@@
For a positive integer $t$, let $tJ=\{tx,tx+1,\ldots,ty-1,ty\}$. First note that every integer $n\in tJ$ has a partition $\lambda$ consisting of $t$ integers in $J$. Indeed, for $n=ty$ we can take $\lambda=(y,\ldots,y)$, and for $tx\leq n\leq ty-1$ we can write $n=tq+r$ for some integers $x\leq q\leq y-1$ and $0\leq r\leq t-1$, and take $\lambda$ to be $r$ $(q+1)$’s followed by (tr)(t-r) $q$’s.
We now see that as $t$ increases, the intervals $tJ$ become closer to each other. Specifically, for $t\geq\left\lceil\frac{x-1}{y-x}\right\rceil\geq\frac{x-1}{y-x}$, we have $ty+1\geq(t+1)x$, which means that as soon as we exit the interval $tJ$, we are already in the interval $(t+1)J$. This means that every integer $n$ satisfying (3.2) is in some interval $tJ$ and has a partition $\lambda$ with all parts in $J$. ∎
@@
Lemma 3.12.
Let $G$, $a$, $b$, $c$, $c_{1}$, and $n$ be as in Notation 3.1. If $n\geq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.